OpenStudy (anonymous):

Whats the volume 10ft by 32ft. cylinder? use pie= 3.14

jimthompson5910 (jim_thompson5910):

what's the height and radius?

jimthompson5910 (jim_thompson5910):

is the height 10? or is the height 32?

OpenStudy (anonymous):

height is 10 radius 32

jimthompson5910 (jim_thompson5910):

ok thanks V = pi*r^2*h V = 3.14*32^2*10 V = ???

OpenStudy (anonymous):

so....whats the answer??

jimthompson5910 (jim_thompson5910):

use a calculator to evaluate 3.14*32^2*10 to get the answer

OpenStudy (anonymous):

how do i do that? sorry im new at this lol

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have

OpenStudy (anonymous):

the normal ones

jimthompson5910 (jim_thompson5910):

you'll have to be more specific

jimthompson5910 (jim_thompson5910):

there is no such thing as "normal"

OpenStudy (anonymous):

the ones with 1-9 and + =

OpenStudy (anonymous):

how u want me to explain it?

jimthompson5910 (jim_thompson5910):

ok idk what you mean, so why not use google as a calculator https://www.google.com/search?hl=&q=3.14*32^2*10&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

jimthompson5910 (jim_thompson5910):

if you type in any math expression, google will act like a calculator and give out the result

OpenStudy (anonymous):

oh ok thank u a WHOLE lol!!!!

jimthompson5910 (jim_thompson5910):

ok no need to be rude

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

i wasnt being rude